OpenStudy (anonymous):

f(x)=x^3. How do I find the domain and range? Please and thank you?

OpenStudy (anonymous):

\[x \in R\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

WELCOME.........

OpenStudy (anonymous):

in this case domain and range both are R

OpenStudy (anonymous):

so if f(x)=x^4 would have the same answer, right?

OpenStudy (anonymous):

IN THIS case domain is R

OpenStudy (anonymous):

\[[0,\infty)\]

OpenStudy (anonymous):

range is \[[0,\infty)\]

OpenStudy (anonymous):

Thank you again. You're so much help.

OpenStudy (anonymous):

any other question.......?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

f(x)= square root x+4-2

OpenStudy (anonymous):

can u please drawn it.......

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it \[\sqrt{x+4}-2\] ??

OpenStudy (anonymous):

\[f(x) equals \sqrt{x+4} -2\]

OpenStudy (anonymous):

here domain is \[x+4\ge 0 \implies x \ge 4\]

OpenStudy (anonymous):

sorry\[x \ge -4\]

OpenStudy (anonymous):

range is R

OpenStudy (anonymous):

Thank you. ^_^